OpenStudy (anonymous):

f(x)= (square root(x^4-16x^2) a. Find the domain of f. b.Find F'(x) c.find te slope of the line normal to the graph of f at x=6

myininaya (myininaya):

domain is when that inside of that radical is zero or positive so you need to solve \[x^4-16x^2 \ge 0\]

OpenStudy (anonymous):

how do i do that?

myininaya (myininaya):

\[x^2(x^2-16) \ge 0\] let \[g(x)=x^2(x^2-16)\] g is zero when x=-4,0,4 So I would draw a number line and plot these numbers and test before and after to see if g>0 -----|-----|------|----- -4 0 4 g(-5) | g(-1)| g(1) | g(5) =+(+)|=+(-)| =+(-)|=+(+) =+ =- =- =+ So the domain is when g=0 or when g>0

OpenStudy (anonymous):

so youwol gtx= 0, 4 ,-4 so yor dmai wod e all numbers such ta x cnnot equal t 0,4,-4?

myininaya (myininaya):

that happens on the intervals and at the numbers: \[(-\infty,-4] U \{0\} U [4,\infty)\]

myininaya (myininaya):

now to find f' you need to use chain rule!

myininaya (myininaya):

\[f'=\frac{1}{2}(x^4-16x^2)^\frac{-1}{2}(x^4-16x^2)'\]

OpenStudy (anonymous):

nd hn he 1/2 is multplid b the (x^4-16x^2)?

myininaya (myininaya):

i don't understand the question

OpenStudy (anonymous):

ok uno how is

myininaya (myininaya):

ok you want to multiply the order is commutative so you can do that

myininaya (myininaya):

\[f'=\frac{2x^3-16x}{\sqrt{x^4-16x^2}} \]